Part A What mass of sodium benzoate should be added to 140.0 mL of a 0.15...

Part A

What mass of sodium benzoate should be added to 140.0 mL of a 0.15 M benzoic acid solution in order to obtain a buffer with a pH of 4.30? Express your answer using two significant figures.
Answer #1

# pH= pkat log (nsalt) nacio) 4-30 = 4-20 + log in saltinacid) 00 0.1-10g (nealt In Acid). 1-2589 = nsalt Nacid nsalt=1-2589if you have any doubts about the answer please mention in the comments please do rate the answer.